# Differential Privacy for Transformer Embeddings with Nonparametric Variational Information Bottleneck

### v1

@@ Abstract
We propose a privacy-preserving method for sharing text data by sharing noisy versions of their transformer embeddings. It has been shown that hidden representations learned by deep models can encode sensitive information from the input, making it possible for adversaries to recover the input data with considerable accuracy. This problem is exacerbated in transformer embeddings because they consist of multiple vectors, one per token. To mitigate this risk, we propose Nonparametric Variational Differential Privacy (NVDP), which ensures both useful data sharing and strong privacy protection. We take a differential privacy approach, integrating a Nonparametric Variational Information Bottleneck (NVIB) layer into the transformer architecture to inject noise into its multi-vector embeddings and thereby hide information, and measuring privacy protection with Rényi divergence and its corresponding Bayesian Differential Privacy (BDP) guarantee. Training the NVIB layer calibrates the noise level according to utility. We test NVDP on the GLUE benchmark and show that varying the noise level gives us a useful tradeoff between privacy and accuracy. With lower noise levels, our model maintains high accuracy while offering strong privacy guarantees, effectively balancing privacy and utility.
@@ 1 Introduction
Deep learning methods are highly dependent on the availability of data, but sharing data is often limited by privacy concerns. This is especially true for text data, where private attributes such as gender or age are mixed with useful information and can be detected by attackers (li-etal-2018-towards). Even when sharing the embeddings of text instead, these representations can still contain sensitive information, and an adversary could use techniques like a GAN (Generative Adversarial Network) attack (hitaj2017deep) to reverse-engineer the original input, potentially reconstructing the text and exposing private information. This is especially true for state-of-the-art attention-based models like transformers, where a text embedding consists of many vectors, one per text token. We want to be able to share such transformer embeddings while still addressing privacy concerns.
@@
Differential Privacy (dwork2006differential) is widely recognized as the benchmark for rigorously quantifying and mitigating privacy risks associated with processing sensitive data. However, integrating differential privacy into machine learning (abadi2016deep; bassily2014private) remains a persistent challenge, primarily due to the substantial drop in model performance compared to non-private versions.
@@
We take the approach of applying differential privacy at the stage of sharing the data, before applying machine learning. This approach has the advantage that the shared data can be reused for multiple purposes and to train multiple models. However, differential privacy relies on noise to remove information, and many types of data, in particular discrete data such as text, do not have a straightforward model of noise. If we first embed the data with a transformer encoder and then share noisy embeddings, then we can apply this approach to any data which can be embedded with a transformer. The transformer architecture has gained prominence for its flexibility and effectiveness across diverse tasks, in particular text.
@@
In this paper, we propose a method for adding noise to transformer embeddings which results in both retaining useful information and differential privacy guarantees. The key to this method is using a nonparametric variational information bottleneck (NVIB) regularizer to train a noise model which is calebrated to the downstream task. Our proposed nonparametric variational differential privacy (NVDP) model first uses NVIB to train a distribution over transformer embeddings, and then samples from this distribution to get a noisy embedding which can be shared. Unlike prior approaches to local differential privacy for sentence embeddings (du2023sanitizing; meehan2022sentence) that rely on task-agnostic noise mechanisms, our NVDP model’s primary novelty lies in its task-aware calibration of randomness. In addition, NVIB calibrates the total randomness across the multiple vectors in a document’s transformer embedding, rather than adding noise independently to individual sentence embeddings. Using Rényi Divergence as a measure of privacy (Geumlek_2017), we show that NVDP provides an effective tradeoff between level of privacy and usefulness in the downstream task.
@@
  • Propose a DP model (NVDP) which uses a VIB regulariser to balance differential privacy against task-aware usefulness.
  • Propose to use a nonparametric VIB regulariser (NVIB) to support sharing multi-vector attention-based representations with differential privacy.
  • Show empirically that NVDP provides a useful tradeoff between privacy and utility.
  • Show empirically that NVIB regularisation is more effective than VIB regularisation used in an analogous way.
@@ 2 Background
Our proposed method builds on previous work on Rényi differential privacy and Nonparametric Variational Information Bottleneck.
@@ 2.1 Differential Privacy
Differential privacy has emerged as the leading standard for rigorously addressing privacy leakage during the processing of sensitive datasets (dwork2014algorithmic). It prevents attackers from deducing too much information about the input data solely from the algorithm’s outputs. [^1] There is no distinction between sensitive information and non-sensitive information, since we can’t tell with certainty what an attacker can deduce from the available information. Thus differential privacy simply tries to reduce all information. There are two primary settings for differential privacy: global differential privacy (GDP) and local differential privacy (LDP). LDP is designed for scenarios where data comes from end users who do not trust a central data collector or third parties with their raw data (dwork2014algorithmic). Under LDP, each user independently perturbs their data before sharing it, ensuring that sensitive information remains protected even if the data collector is compromised. By introducing noise or randomness at the user level, LDP enables the collection of aggregated insights from a dataset without revealing individual-level details, thus maintaining privacy while still allowing for statistical analysis (dwork2014algorithmic; abadi2016deep). In our work, we aim to protect embeddings of text by sharing noisy embeddings, and thus we apply local differential privacy. To formalize this, differential privacy ensures that for every pair of adjacent inputs $\bm{x}$} and $\bm{x}$$^{\prime}$ and a randomized algorithm $M$, the distribution of $M({\bm{x}})$ and $M({\bm{x}}^{\prime})$ are close to each other. Originally, the distributions of $M({\bm{x}})$ and $M({\bm{x}}^{\prime})$ are considered adjacent if they differ by all the attributes of a single record, but in practice the notion of adjacency can vary. For example, sun-etal-2019-mitigating define two sentences as adjacent if they differ by at most 5 consecutive words, where each sentence is split into segments of 5 words. In our framework, we operate under the standard Local Differential Privacy (LDP) model, where the guarantee concerns the indistinguishability of the mechanism’s output for any two distinct inputs $\bm{x}$} and $\bm{x}$$^{\prime}$. Thus, we consider all inputs $\bm{x}$} to be neighbors of all other inputs $\bm{x}$$^{\prime}$.
@@
is a variation of standard $(\epsilon,\delta)$-differential privacy that employs Rényi divergence as a metric to measure the distance between the output distributions $Q\text{ of }M({\bm{x}})$ and $Q^{\prime}\text{ of }M({\bm{x}}^{\prime})$ (Geumlek_2017). This approach is especially advantageous for training machine learning models that require differential privacy.
Definition 2.1.
Given two probability distributions $\mathcal{Q}$ and $\mathcal{Q^{\prime}}$ defined over domain $\mathcal{Z}$, the Rényi divergence of order $\lambda>1$ is:
$\hskip-25.83325ptD_{\lambda}(Q||Q^{\prime})=\frac{1}{\lambda-1}\log\left(\int_{\bm{z}}Q({\bm{z}})\left(\frac{Q({\bm{z}})}{Q^{\prime}({\bm{z}})}\right)^{\lambda-1}\right)\vskip-4.30554pt$ (1)
@@
A randomized mechanism $M:\mathcal{X}\rightarrow\mathcal{Z}$ satisfies $(\lambda,\epsilon)$-Local Rényi Differential Privacy if for any pair of adjacent inputs ${\bm{x}},{\bm{x}}^{\prime}\in\mathcal{X}$, the following holds:
$D_{\lambda}(M({\bm{x}})||M({\bm{x}}^{\prime}))\leq\epsilon$ (2)
The privacy parameter $\epsilon$, often called privacy budget, measures the privacy loss associated by the output of the algorithm: $\epsilon=0$ provides perfect privacy, meaning the output is completely independent of the input. As $\epsilon\rightarrow\inf$, the privacy guarantees diminish, offering no protection for the input data. The other parameter $\lambda$ controls the importance of the worst-case privacy violations, reducing to KL-divergence at its lowest value and the maximum log-difference at its highest value.
@@
offers an alternative interpretation of privacy by focusing on the change in an adversary’s belief. As introduced by triastcyn2020bayesian, instead of comparing two adjacent inputs $\bm{x}$} and $\bm{x}$$^{\prime}$, BDP analyzes how an output from a mechanism $M({\bm{x}})$ changes an adversary’s posterior belief about the input $\bm{x}$} compared to their prior belief. This prior is typically formed from knowledge of the underlying data distribution, which we can denote as $\mathcal{X}$. Formally, this is captured in the $(\epsilon_{\mu},\delta_{\mu})\text{-BDP}$ definition:
Definition 2.3.
Let $M:\mathcal{X}\rightarrow\mathcal{Z}$ be a randomized mechanism applied to individual data points. Then $M$ satisfies $(\epsilon_{\mu},\delta_{\mu})$-Local Bayesian Differential Privacy if, for all $\bm{x}$}, a data point $\bm{x}$$^{\prime}$ drawn from the data distribution ${\bm{x}}^{\prime}\sim\mathcal{X}$, and all measurable subsets $S\subseteq\mathcal{Z}$, the following holds:
$\Pr[M({\bm{x}})\in S]\leq e^{\epsilon_{\mu}}\Pr[M({\bm{x}}^{\prime})\in S]+\delta_{\mu}\vskip-4.30554pt$ (3)
Here, $\Pr[M({\bm{x}}^{\prime})\in S]$ marginalizes over both the outputs in $S$ and the datapoints ${\bm{x}}^{\prime}\sim\mathcal{X}$, and $\delta_{\mu}$ accounts not only for the probability of catastrophic privacy loss from the mechanism itself but also incorporates the uncertainty an analyst has about the true data distribution. This framework provides a practical way to reason about privacy for typical data points, which is highly relevant for machine learning models trained on specific data domains.
@@
triastcyn2020bayesian also point out that Rényi divergence can be used to define an upper bound on the worst-case privacy loss over sampled outputs. They use this fact to define a privacy accountant which allows multiple accesses to the data, but it can also be used to simply define a new privacy measure which is less sensitive to worst-case loss and closer to expected-case loss over outputs, depending on the setting of $\lambda$. This gives us a privacy measure which summarises the privacy risk over both the distribution of alternative examples $x^{\prime}$ and the possible outputs. In our experiments, we leverage this connection by first calculating the Rényi Divergence (RD) and then converting it into an interpretable $(\epsilon_{\mu},\delta_{\mu})$-BDP guarantee. We use the privacy accounting mechanism detailed in Theorem $2$ of triastcyn2020bayesian. For the full derivation and implementation details of the accountant, we refer the reader to the original work by triastcyn2020bayesian.
@@ 2.2 Nonparametric Variational Information Bottleneck
NVIB (hendersonfehr2023) is a variational information bottleneck regulariser for attention layers. It replaces the set of key vectors in a transformer’s attention layer with a latent mixture of impulse distributions, specified as a set of vectors $\bm{Z}$ and their normalised weights $\bm{\pi}$. To access this mixture distribution, it generalises the attention function to “denoising attention”. It then uses Bayesian nonparametrics to define prior and posterior distributions over these mixture distributions, and a KL divergence with the prior to regularise the amount of information in its posterior.
@@
Since NVIB uses Bayesian inference to specify its posterior, it needs a prior over the latent space. Since the number of keys in an attention layer grows with the length of the input and the input can be arbitrarily large, this prior needs to specify a distribution over sets of weighted vectors $(\bm{\pi}\in\mathbb{R}^{m},\bm{Z}\in\mathbb{R}^{m\times d})$ which have no finite limit to their size $m$. Nevertheless, we can still define probability distributions over this infinite set by applying Bayesian nonparametric methods. In particular, hendersonfehr2023 use a Dirichlet Process (DP) to define the prior distribution $DP(G^{p}_{0},\alpha^{p}_{0})$ over unboundedly large mixtures (“p” designates the prior and “q” designates the posterior). The vectors $\bm{Z}_{i}$ are generated by the base distribution $G^{p}_{0}$, which is a Guassain distribution with parameters $(\bm{\mu^{p}}=\bm{0},(\bm{\sigma^{p}})^{2}=\bm{1})$. The weights $\bm{\pi}$ are generated with a symmetric Dirichlet distribution parameterised by the total pseudo-count $\alpha_{0}^{p}=1$, which tends to generate large weights on a few vectors and a long tail of exponentially smaller weights.
@@
An NVIB layer uses Bayesian inference to parameterise its posterior distribution $DP(G^{q}_{0},\alpha^{q}_{0})$ in terms of a set of pseudo-observations computed from the set of vectors which the transformer inputs to the layer. These $n$ input vectors $\bm{x}\in\mathbb{R}^{n\times d}$ are each individually projected to a pseudo-count $\alpha^{q}_{i}$$\in\mathbb{R}$ and a pair of Gaussian parameters $(\bm{\mu^{q}_{i}}\in\mathbb{R}^{d},(\bm{\sigma^{q}_{i}})^{2}\in\mathbb{R}^{d})$. The base distribution $G^{q}_{0}$ is a mixture of Gaussians, where each pair of Gaussian parameters specifies a component of the mixture, and the pseudo-counts specify their weights. In addition, there is an $n{+}1^{\text{th}}$ component of the base distribution specified by the prior’s parameters, so $(\alpha^{q}_{n+1},\bm{\mu^{q}_{n+1}},(\bm{\sigma^{q}_{n+1}})^{2})=(\alpha^{p}_{0},\bm{\mu^{p}},(\bm{\sigma^{p}})^{2})$.
$F$ $\sim\operatorname*{DP}\left(G^{q}_{0},~\alpha^{q}_{0}\right)$ $\alpha^{q}_{0}$ $=\sum_{i=1}^{n+1}{\alpha}^{q}_{i}$ $G^{q}_{0}$ $=\sum_{i=1}^{n+1}\frac{{\alpha}^{q}_{i}}{\alpha^{q}_{0}}{G}^{q}_{i}$ $G^{q}_{i}$ $=\mathcal{N}(\bm{\mu^{q}_{i}},\bm{I}(\bm{\sigma^{q}_{i}})^{2})\vskip-4.30554pt$ (4)
As in previous work (fehr-colm2024-nonparametric), during training our NVIB layer approximates samples of weighted vectors from this posterior as $\bm{\pi}\sim\text{Dir}(\bm{\alpha^{q}})$ and $\bm{Z_{i}}\sim\mathcal{N}(\bm{\mu^{q}_{i}},\bm{(\sigma^{q}_{i})}^{2})$, where $Dir(\cdot)$ is a Dirichlet distribution. An NVIB layer also has the capability to drop specific embeddings by setting their pseudo-counts to zero, effectively reducing the complexity of the representation.
@@
The NVIB loss regularises the flow of information through the latent representation. The loss consists of three components: a task loss $(L_{T})$ and two Kullback-Leibler (KL) divergence terms, $L_{D}$ and $L_{G}$. The $L_{T}$ term serves as a supervised learning objective, guiding the latent representation to retain enough information to perform the task. The $L_{G}$ term encourages noise in the Gaussian components, making them each less informative. The $L_{D}$ term both encourages noise in the weights and encourages some of the Dirichlet distribution’s pseudo-counts to reach zero, effectively eliminating some vectors and reducing the overall information capacity of the latent representation. Two hyperparameters are added to adjust the impact of both the $L_{D}$ and $L_{G}$ terms:
$L=L_{T}+\lambda_{D}L_{D}+\lambda_{G}L_{G}\vskip-2.15277pt$ (5)

### v2

@@ Abstract
We propose a privacy-preserving method for sharing text data by sharing noisy versions of their transformer embeddings. It has been shown that hidden representations learned by deep models can encode sensitive information from the input, making it possible for adversaries to recover the input data with considerable accuracy. This problem is exacerbated in transformer embeddings because they consist of multiple vectors, one per token. To mitigate this risk, we propose Nonparametric Variational Differential Privacy (NVDP), which ensures both useful data sharing and strong privacy protection. We take a differential privacy (DP) approach, integrating a nonparametric variational information bottleneck (NVIB) layer into the transformer architecture to inject noise into its multivector embeddings and thereby hide information, and measuring privacy protection with Rényi Divergence (RD) and its corresponding Bayesian Differential Privacy (BDP) guarantee. Training the NVIB layer calibrates the noise level according to the utility of the downstream task. We test NVDP on the General Language Understanding Evaluation (GLUE) benchmark and show that varying the noise level gives us a useful trade-off between privacy and accuracy. With lower noise levels, our model maintains high accuracy while offering strong privacy guarantees, effectively balancing privacy and utility.
@@ 1 Introduction
Deep learning methods are highly dependent on the availability of data, but sharing data is often limited by privacy concerns. This is especially true for text data, where private attributes such as gender or age are mixed with useful information and can be detected by attackers (Li et al., 2018). Even when sharing the embeddings of text instead, these representations can still contain sensitive information, and an adversary could use techniques like a GAN (Generative Adversarial Network) attack (Hitaj et al., 2017) to reverse-engineer the original input, potentially reconstructing the text and exposing private information. This is especially true for state-of-the-art attention-based models like transformers, where a text embedding consists of many vectors, one per text token. We want to be able to share such transformer embeddings while still addressing privacy concerns.
@@
Differential Privacy (DP) (Dwork, 2006) is widely recognized as the benchmark for rigorously quantifying and mitigating privacy risks associated with processing sensitive data. However, integrating DP into machine learning (Abadi et al., 2016; Bassily et al., 2014) remains a persistent challenge, primarily due to the substantial drop in model performance compared to non-private versions.
@@
We take the approach of applying DP at the stage of sharing the data, before applying machine learning. This approach has the advantage that the shared data can be reused for multiple purposes and to train multiple models. However, DP relies on noise to remove information, and many types of data, in particular discrete data such as text, do not have a straightforward model of noise. If we first embed the data with a transformer encoder and then share noisy embeddings, then we can apply this approach to any data which can be embedded with a transformer. The transformer architecture has gained prominence for its flexibility and effectiveness across diverse tasks, in particular text.
@@
In this paper, we propose a method for adding noise to transformer embeddings which results in both retaining useful information and DP guarantees. The key to this method is using a nonparametric variational information bottleneck (NVIB) regularizer to train a noise model which is calibrated to the downstream task. Our proposed nonparametric variational differential privacy (NVDP) model first uses NVIB to train a distribution over transformer embeddings, and then samples from this distribution to get a noisy embedding which can be shared. Unlike prior approaches to local DP for sentence embeddings (Du et al., 2023; Meehan et al., 2022) that rely on task-agnostic noise mechanisms, our NVDP model’s primary novelty lies in its task-aware calibration of randomness. In addition, NVIB calibrates the total randomness across the multiple vectors in a document’s transformer embedding, rather than adding noise independently to individual sentence embeddings. Using Rényi Divergence (RD) as a measure of privacy (Geumlek et al., 2017), we show that NVDP provides an effective trade-off between level of privacy and usefulness in the downstream task.
@@
  • Propose a DP model (NVDP) which uses a variational information bottleneck (VIB) regularizer to balance DP against task-aware usefulness.
  • Propose to use a nonparametric VIB regularizer (NVIB) to support sharing multivector attention-based representations with DP.
  • Show empirically that NVDP provides a useful trade-off between privacy and utility.
  • Show empirically that NVIB regularization is more effective than VIB regularization used in an analogous way.
@@ 2 Background
Our proposed method builds on previous work on Rényi differential privacy and NVIB.
@@ 2.1 Differential Privacy
DP has emerged as the leading standard for rigorously addressing privacy leakage during the processing of sensitive datasets (Dwork et al., 2014). It prevents attackers from deducing too much information about the input data solely from the algorithm’s outputs. [^1] There is no distinction between sensitive information and non-sensitive information, since we can’t tell with certainty what an attacker can deduce from the available information. Thus DP simply tries to reduce all information. There are two primary settings for DP: global differential privacy (GDP) and local differential privacy (LDP). LDP is designed for scenarios where data comes from end users who do not trust a central data collector or third parties with their raw data (Dwork et al., 2014). Under LDP, each user independently perturbs their data before sharing it, ensuring that sensitive information remains protected even if the data collector is compromised. By introducing noise or randomness at the user level, LDP enables the collection of aggregated insights from a dataset without revealing individual-level details, thus maintaining privacy while still allowing for statistical analysis (Dwork et al., 2014; Abadi et al., 2016). In our work, we aim to protect embeddings of text by sharing noisy embeddings, and thus we apply LDP. To formalize this, DP ensures that for every pair of adjacent inputs $\bm{x}$} and $\bm{x}$$^{\prime}$ and a randomized algorithm $M$, the distribution of $M({\bm{x}})$ and $M({\bm{x}}^{\prime})$ is close to each other. Originally, the distributions of $M({\bm{x}})$ and $M({\bm{x}}^{\prime})$ are considered adjacent if they differ by all the attributes of a single record, but in practice the notion of adjacency can vary. For example, Sun et al. (2019) defines two sentences as adjacent if they differ by at most 5 consecutive words, where each sentence is split into segments of 5 words. In our framework, we operate under the standard LDP model, where the guarantee concerns the indistinguishability of the mechanism’s output for any two distinct inputs $\bm{x}$} and $\bm{x}$$^{\prime}$. Thus, we consider all inputs $\bm{x}$} to be neighbors of all other inputs $\bm{x}$$^{\prime}$.
@@
is a variation of standard $(\epsilon,\delta)$-DP that employs RD as a metric to measure the distance between the output distributions $Q\text{ of }M({\bm{x}})$ and $Q^{\prime}\text{ of }M({\bm{x}}^{\prime})$. This approach is especially advantageous for training machine learning models that require DP (Mironov, 2017).
Definition 2.1 (Rényi Divergence).
Given two probability distributions $Q$ and $Q^{\prime}$ defined over domain $\mathcal{Z}$, the RD of order $\lambda>1$ is:
$\hskip-25.83325ptD_{\lambda}(Q||Q^{\prime})=\frac{1}{\lambda-1}\log\left(\int_{\bm{z}}Q({\bm{z}})\left(\frac{Q({\bm{z}})}{Q^{\prime}({\bm{z}})}\right)^{\lambda-1}\right).\vskip-4.30554pt$ (1)
@@
A randomized mechanism $M:\mathcal{X}\rightarrow\mathcal{Z}$ satisfies $(\lambda,\epsilon)$-RDP if for any pair of adjacent inputs ${\bm{x}},{\bm{x}}^{\prime}\in\mathcal{X}$, the Rényi divergence of order $\lambda$ between their output distributions is bounded by $\epsilon$:
$D_{\lambda}(M({\bm{x}})||M({\bm{x}}^{\prime}))\leq\epsilon.$ (2)
The privacy parameter $\epsilon$, often called privacy budget, measures the privacy loss associated by the output of the algorithm: $\epsilon=0$ provides perfect privacy, meaning the output is completely independent of the input. As $\epsilon\rightarrow\inf$, the privacy guarantees diminish, offering no protection for the input data. The other parameter $\lambda$ controls the importance of the worst-case privacy violations, reducing to Kullback–Leibler (KL) divergence at its lowest value and the maximum log-difference at its highest value.
@@
offers an alternative interpretation of privacy by focusing on the change in an adversary’s belief. As introduced by Triastcyn and Faltings (2020), instead of comparing two adjacent inputs $\bm{x}$} and $\bm{x}$$^{\prime}$, BDP analyzes how an output from a mechanism $M({\bm{x}})$ changes an adversary’s posterior belief about the input $\bm{x}$} compared to their prior belief. This prior is typically formed from knowledge of the underlying data distribution, which we can denote as $\mathcal{X}$. Formally, this is captured in the $(\epsilon_{\mu},\delta_{\mu})\text{-BDP}$ definition:
Definition 2.3 ($(\epsilon_{\mu},\delta_{\mu})$-Bayesian Differential Privacy).
A randomized mechanism $M:\mathcal{X}\rightarrow\mathcal{Z}$ satisfies $(\epsilon_{\mu},\delta_{\mu})$-BDP if for any individual data point ${\bm{x}}\in\mathcal{X}$ and all measurable subsets $S\subseteq\mathcal{Z}$:
$\Pr[M({\bm{x}})\in S]\leq e^{\epsilon_{\mu}}\Pr[M({\bm{x}}^{\prime})\in S]+\delta_{\mu},$ (3)
where $\bm{x}$$^{\prime}$ is a data point sampled from the underlying data distribution $\mathcal{X}$.
In this framework, $\Pr[M({\bm{x}}^{\prime})\in S]$ represents the marginal probability of an output falling in $S$ over the data distribution $\mathcal{X}$. The parameter $\delta_{\mu}$ not only accounts for the probability of catastrophic privacy loss from the mechanism itself, but also incorporates the uncertainty regarding the true data distribution (Triastcyn and Faltings, 2020). This provides a pragmatic way to reason about privacy for typical data points, making it highly relevant for models trained on domain-specific datasets.
@@
Triastcyn and Faltings (2020) also points out that RD can be used to define an upper bound on the worst-case privacy loss over sampled outputs. They use this fact to define a privacy accountant which allows multiple accesses to the data, but it can also be used to simply define a new privacy measure which is less sensitive to worst-case loss and closer to expected case loss over outputs, depending on the setting of $\lambda$. This gives us a privacy measure which summarises the privacy risk over both the distribution of alternative examples $x^{\prime}$ and the possible outputs. In our experiments, we leverage this connection by first calculating the RD and then converting it into an interpretable $(\epsilon_{\mu},\delta_{\mu})$-BDP guarantee. We use the privacy accounting mechanism detailed in Theorem $2$ of Triastcyn and Faltings (2020). For the full derivation and implementation details of the accountant, we refer the reader to the original work by Triastcyn and Faltings (2020).
@@ 2.2 Nonparametric Variational Information Bottleneck
NVIB (Henderson and Fehr, 2023) is a VIB regularizer for attention layers. It replaces the set of key vectors in a transformer’s attention layer with a latent mixture of impulse distributions, specified as a set of vectors $\bm{Z}$ and their normalized weights $\bm{\pi}$. To access this mixture distribution, it generalizes the attention function to “denoising attention”. It then uses Bayesian nonparametrics to define prior and posterior distributions over these mixture distributions, and a KL divergence with the prior to regularize the amount of information in its posterior.
@@
Since NVIB uses Bayesian inference to specify its posterior, it needs a prior over the latent space. Since the number of keys in an attention layer grows with the length of the input and the input can be arbitrarily large, this prior needs to specify a distribution over sets of weighted vectors $(\bm{\pi}\in\mathbb{R}^{m},\bm{Z}\in\mathbb{R}^{m\times d})$ which have no finite limit to their size $m$. Nevertheless, we can still define probability distributions over this infinite set by applying Bayesian nonparametric methods. In particular, Henderson and Fehr (2023) use a Dirichlet Process to define the prior distribution $DP(G^{p}_{0},\alpha^{p}_{0})$ over unboundedly large mixtures (“p” designates the prior and “q” designates the posterior). The vectors $\bm{Z}_{i}$ are generated by the base distribution $G^{p}_{0}$, which is a Gaussian distribution with parameters $(\bm{\mu}^{p}=\bm{0},(\bm{\sigma}^{p})^{2}=\bm{1})$. The weights $\bm{\pi}$ are generated with a symmetric Dirichlet distribution parameterized by the total pseudo-count $\alpha_{0}^{p}=1$, which tends to generate large weights on a few vectors and a long tail of exponentially smaller weights.
@@
An NVIB layer uses Bayesian inference to parameterize its posterior distribution $DP(G^{q}_{0},\alpha^{q}_{0})$ in terms of a set of pseudo-observations computed from the set of vectors which the transformer inputs to the layer. These $n$ input vectors $\bm{x}\in\mathbb{R}^{n\times d}$ are each individually projected to a pseudo-count $\alpha^{q}_{i}$$\in\mathbb{R}$ and a pair of Gaussian parameters $(\bm{\mu}^{q}_{i}\in\mathbb{R}^{d},(\bm{\sigma}^{q}_{i})^{2}\in\mathbb{R}^{d})$. The base distribution $G^{q}_{0}$ is a mixture of Gaussians, where each pair of Gaussian parameters specifies a component of the mixture, and the pseudo-counts specify their weights. In addition, there is an $n{+}1^{\text{th}}$ component of the base distribution specified by the prior’s parameters, so $(\alpha^{q}_{n+1},\bm{\mu}^{q}_{n+1},(\bm{\sigma}^{q}_{n+1})^{2})=(\alpha^{p}_{0},\bm{\mu}^{p},(\bm{\sigma}^{p})^{2})$.
$F$ $\sim\operatorname*{DP}\left(G^{q}_{0},~\alpha^{q}_{0}\right),$ $\alpha^{q}_{0}$ $=\sum_{i=1}^{n+1}{\alpha}^{q}_{i},$ $G^{q}_{0}$ $=\sum_{i=1}^{n+1}\frac{{\alpha}^{q}_{i}}{\alpha^{q}_{0}}{G}^{q}_{i},$ $G^{q}_{i}$ $=\mathcal{N}(\bm{\mu}^{q}_{i},\bm{I}(\bm{\sigma}^{q}_{i})^{2}).\vskip-4.30554pt$ (4)
As in previous work (Fehr and Henderson, 2024), during training our NVIB layer approximates samples of weighted vectors from this posterior as $\bm{\pi}\sim\text{Dir}(\bm{\alpha}^{q})$ and $\bm{Z_{i}}\sim\mathcal{N}(\bm{\mu}^{q}_{i},(\bm{\sigma}^{q}_{i})^{2})$, where $Dir(\cdot)$ is a Dirichlet distribution. An NVIB layer also has the capability to drop specific embeddings by setting their pseudo-counts to zero, effectively reducing the complexity of the representation.
@@
The NVIB loss regularizes the flow of information through the latent representation. The loss consists of three components: a task loss $(L_{T})$ and two KL divergence terms, $L_{D}$ and $L_{G}$. The $L_{T}$ term serves as a supervised learning objective, guiding the latent representation to retain enough information to perform the task. The $L_{G}$ term encourages noise in the Gaussian components, making them each less informative. The $L_{D}$ term both encourages noise in the weights and encourages some of the Dirichlet distribution’s pseudo-counts to reach zero, effectively eliminating some vectors and reducing the overall information capacity of the latent representation. Two hyperparameters are added to adjust the impact of both the $L_{D}$ and $L_{G}$ terms:
$L=L_{T}+\lambda_{D}L_{D}+\lambda_{G}L_{G}.\vskip-2.15277pt$ (5)